Do you expect the detonating reaction between hydrogen gas and oxygen gas forming water vapor (oxyhydrogen reaction) to have a temperature limit above or below which the reaction will not be spontaneous? Explain your answer and estimate such temperature from thermodynamic data found online.2. (2 pts) Do you expect the detonating reaction between hydrogen gas and oxygen gas forming water vapor (oxyhydrogen reactio
Answer #1

2H, (१) + 2) >24,0(2) ACKJ/mol) As° I/k.mo 130.7 | १) २०.2 H50 (3) 1868 step -241-8 AH = AH°products - 4H reactant = 2241.6-(AG=AH-TAS AS = As products - 45 reactanty = (2x188.8) - ( 2 (130.7) + 205.2) = -89 J/k = -0.089 KJ/K if AG=0 AHETAS - -483.6kAbove the temperature of 5434K, TAS term becomes more than AH and reaction becomes non-spontaneay (since AG becomes positive)